# Floer-theoretic filtration on Painlevé Hitchin systems

### v1

@@ 3.4. Painlevé IV (degenerate) case
With the above assumptions, the trace of $\theta$ vanishes and its determinant reads as
@@
This follows from [ISS19a, Equation (70)]. Note that in that formula, $z_{2}=z^{-1}$, and we made use of the trivialization $\frac{\operatorname{d}\!z_{2}}{z_{2}}$ of $K(D)$ near $z=\infty$. Furthermore, the parameter $b_{-1}$ vanishes and our free parameter $b$ can be identified with the parameter $t$ appearing in [ISS19a]. Finally, $w_{2}=\zeta$. ∎
@@
There exists a $I$-holomorphic $\mathbb{C}^{*}$-action on $\mathcal{M}^{PIV}$, equivariant with respect to the action $b\mapsto t^{3}b$ on $\mathcal{B}^{PIV}$. It acts on $\Omega_{I}^{PIV}$ with weight $2$.
@@
in particular $F_{\mathrm{min}}=F_{0}$. Therefore, subspaces of $H_{1}$ corresponds to flow within $\mathbb{C}P^{1}$’s, thus the weight decomposition of $F_{1},F_{2}$ and $F_{3}$ has $H_{-1},$ thus $H_{3}$ weight space:
(30) $T_{F_{1}}Y=H_{-1}\oplus H_{3},\quad T_{F_{2}}Y=H_{-1}\oplus H_{3},\quad T_{F_{3}}Y=H_{-1}\oplus H_{3}.$
and $H_{3}$ in (30) go outside of the core. As before, the $\mathbb{C}^{*}$-flowline corresponding to $H_{3}$ from one of these fixed points is the Hitchin section. Now, computation of the indices (11) gives:131313where, as before we denote $\mu_{\lambda+}:=\mu_{\lambda+\delta}$ for $01\gg\delta>$ 0
@@
The column $H^{*}(B_{1})$ is determined by 2.10, as the intersection form of $\mathcal{M}^{PIV}$ has a one-dimensional kernel, spanned by the sum of core components (labelled by $E_{1,2,3}$):
@@ 3.5. Painlevé VI case
the $\mathbb{C}^{*}$-fixed locus has to have a total cohomology rank of 6. The $\mathbb{C}P^{1}$ corresponding to the central vertex has 4 fixed points, hence is fixed. Due to the smoothness of fixed loci, no other $\mathbb{C}P^{1}$s are fixed, hence they have another fixed point each, thus:
@@
and $H^{0}(F_{i})[-2]=\langle[E_{i}]\rangle$ in $H^{2}(Y)$, where $E_{i}$ is the $\mathbb{C}P^{1}$ which $F_{i}$ belongs to. By 2.6, $1\in H^{0}(F_{0})$ enters the filtration at $\lambda=2,$ hence altogether:
(34) $\begin{split}\mathscr{F}_{1/2}\cong\mathbb{K}^{4}\subset\mathscr{F}_{1}=H^{2}(Y)\subset\mathscr{F}_{2}=H^{*}(Y)\\ \mathscr{F}_{\mathbb{B},1/2}=\langle[E_{1}],\dots,[E_{4}]\rangle\subset\mathscr{F}_{\mathbb{B},1}=H^{2}(Y)\subset\mathscr{F}_{2}=H^{*}(Y)\end{split}$
From the weight decompositions (32),(33) we get the outer torsion manifolds $Y_{2}\cong\mathbb{C}\sqcup\mathbb{C}\sqcup\mathbb{C}\sqcup\mathbb{C},$ thus $B_{1/2}\cong S^{1}\sqcup S^{1}\sqcup S^{1}\sqcup S^{1}.$ Their shifts, computed via (15), give the columns of the $E_{1}$-page of spectral sequence, shown on Figure 4. The column $H^{*}(B_{1})$ is determined by 2.10, as the intersection form of $\mathcal{M}^{PVI}$ has a one-dimensional kernel, spanned by the weighted sum $2[F_{0}]+[C_{i}]$ of the core components. Like in the previous example, the spectral sequence alone does not completely determine the filtration (34), as we could not tell which of the columns $B_{1},B_{3/2}$ and $B_{2}$ annihilates $1\in H^{0}(\mathcal{M}^{PIV})$.
@@ 3.6. Remaining Painlevé cases
3.7. Other Higgs moduli of $\dim_{\mathbb{C}}=2$
@@
induces Hitchin fibration
(36) $h_{\Gamma}:\mathcal{M}_{\Gamma}\rightarrow\mathbb{C}/\Gamma=\mathcal{B}_{\Gamma}\cong\mathbb{C}.$
@@
[Zha17, Proposition 5.4] The central fibre of (36) is
@@
The fiber-contraction on $T^{*}E$ that makes (35) equivariant (with weight-1 action on $\mathbb{C}$) induces a $\mathbb{C}^{*}$-action on $\mathcal{M}_{\Gamma}$ that makes (36) equivariant. For completeness, we describe this action in the Higgs bundle-theoretic way:
@@
$\zeta^{4}+b\frac{(\operatorname{d}\!z)^{\otimes 4}}{z^{3}(z-1)^{3}}.$
@@
$\zeta^{6}+b\frac{(\operatorname{d}\!z)^{\otimes 6}}{z^{5}(z-1)^{4}},$
@@
The Floer-theoretic filtration $\mathscr{F}$ with respect to this $\mathbb{C}^{*}$-action on $\mathcal{M}_{\Gamma}$ is studied in [RŽ23b], where it is compared with the perverse filtration of Hitchin fibration $h_{\Gamma}:\mathcal{M}_{\Gamma}\rightarrow\mathcal{B}_{\Gamma}$ (the famous “$P=W$” filtration of Higgs moduli spaces). In addition, there is an interesting correlation with the associated root system of graph $Q_{\Gamma}.$ Being an affine Dynkin graph, it has the imaginary root, the positive $\mathbb{Z}$-generator of the radical of the symmetric bilinear form with associated Cartan matrix $C_{Q_{\Gamma}}.$ One can label the vertices $i\in Q_{\Gamma}^{0}$ by the corresponding entries $n_{i}$ of the imaginary root, see Figure 5
@@
Given two filtrations $F_{i}$ and $G_{j}$ of a vector space $V,$ we say that $F$ refines $G$ if for each $j$ there is some $i$ such that $G_{j}=F_{i}$.
@@ 4. Three filtrations compared
The imaginary root labels mentioned above have another geometric meaning. Recall from Poposition 3.11 that $\mathrm{Core}(\mathcal{M}_{\Gamma})=h_{\Gamma}^{-1}(0)=\cup E_{i}$ has irreducible components $E_{i}$ labelled by vertices $i$ of $Q_{\Gamma}.$ Moreover, using its canonical scheme structure, it comes with multiplicities $m_{i}$ of curves $E_{i}$
@@
with the labellings $n_{i}$ of the imaginary root of $Q_{\Gamma}.$ This is a consequence of
@@
    Self-intersection of $h_{\Gamma}^{-1}(0)$ (with multiplicities) in the sublattice spanned by $\mathrm{Irr}(\mathrm{Core}(\mathcal{M}))$ is equal to zero, as a fibre class
@@
    by definition, the imaginary root is the positive integer generator of the radical of the symmetric bilinear form with given Cartan matrix, which equals minus the intersection matrix of the components $E_{i}$ of $\mathrm{Irr}(\mathrm{Core}(\mathcal{M}))$.
@@
$\mathrm{rk}\,\mathscr{F}_{k/|\Gamma|}=\#\{E_{i}\in\mathrm{Irr}(\mathrm{Core}(\mathcal{M}_{\Gamma}))\mid m_{i}\leq k\}$
@@
for any 2-dimensional parabolic Higgs bundle $\mathcal{M}_{\Gamma}.$ Following ideas from [RŽ23a], one expects that these filtrations actually coincide, not only rank-wise:
@@
Given a $2$-dimensional Higgs moduli space $\mathcal{M}$ which has an equivariant $I$-holomorphic $\mathbb{C}^{*}$-action and such that the Hitchin map has generically smooth fibers, its multiplicity and Floer-theoretic filtration coincide rank-wise. For moduli spaces of parabolic Higgs bundles they both refine the $P=W$ filtration, whereas for Painlevé spaces, $P=W$ refines the former two filtrations. Thus, in the intersection, i.e. for Painlevé VI space, these three filtrations coincide.
@@
We first claim that a Higgs moduli space $\mathcal{M}$ satisfying the conditions is diffeomorphic to either of Painlevé type or one of 5 parabolic $\mathcal{M}_{\Gamma}$. Indeed, let $\mathcal{M}$ be a $2$-dimensional Higgs moduli space. Assume that the generic fiber of the Hitchin map $h$ is smooth. Then, $\mathcal{M}$ is an elliptic fibration over $\mathbb{C}$. Now, assume that $\mathcal{M}$ carries a $\mathbb{C}^{*}$-action for which $h$ is equivariant. Clearly, then $h$ has at most one singular fiber, namely over 00. If $h$ admits no singular fiber, then $\mathcal{M}$ is diffeomorphic to $T^{*}E$ for an elliptic curve $E$, i.e. $\mathcal{M}_{\Gamma}$ for $\Gamma=0$. Assume that $h$ admits exactly one singular fiber. It is known that $\mathcal{M}$ is a smooth quasi-projective surface and $h$ is a dominant algebraic morphism to $\mathbb{C}$. Therefore, there exists a smooth projective surface $\mathcal{M}\subset\overline{\mathcal{M}}$ such that $\overline{\mathcal{M}}\setminus\mathcal{M}$ is a simple normal crossing divisor and a morphism
$\overline{h}\colon\overline{\mathcal{M}}\to\mathbb{C}P^{1}$
such that $\overline{h}|_{\mathcal{M}}=h$. Now, it is known that if a smooth elliptic fibration over $\mathbb{C}P^{1}$ has exactly two singular fibers, then these two singular fibers must be one of the possibilities
$\widetilde{E}_{8}+II,\qquad\widetilde{E}_{7}+III,\qquad\widetilde{E}_{6}+IV,\qquad\widetilde{D}_{4}+\widetilde{D}_{4}.$
In each of the cases (except for the last one), $\overline{h}^{-1}(0)$ may be either the first of the two singular curves listed and $\overline{h}^{-1}(\infty)$ the second one, or vice versa. These combinations give rise to $\mathcal{M}_{\Gamma}$ with $\Gamma\neq 0$ and to the Painlevé I, II, IV, VI cases, respectively. As we have seen in Proposition 3.10, amongst the Painlevé spaces, only I, II, IV and VI admit $\mathbb{C}^{*}$-actions. The intersection of parabolic and Painlevé spaces is precisely the Painlevé VI case, i.e. $\mathcal{M}_{\mathbb{Z}/2}$ space, whose core consists of an affine Dynkin $\widetilde{D}_{4}$ tree of rational curves.
By (38), the multiplicity and Floer-theoretic filtration coincide rank-wise for parabolic spaces. For Painlevé I, II, IV spaces, both filtrations are trivial, so they coincide. The Floer-theoretic filtration $\mathscr{F}=\mathscr{F}_{\mathbb{B}}(H^{2}(\mathcal{M}))$ is trivial due to (20),(27),(31), respectively. The multiplicity filtration is trivial as the multiplicities of core components $E_{i}$ are all $m_{i}=1$ in these cases. The reasoning is similar as for (37), as the self-intersection of $\sum_{i}E_{i}$ is zero. The latter is already proved at the ends of their respective Sections 3.1, 3.3 and 3.4.
For the parabolic moduli spaces $\mathcal{M}_{\Gamma}$, the multiplicity filtration $\mathscr{M}$ refines the $P=W$, as for $\Gamma\neq 0,$ the latter has two steps:
@@
For Painlevé I, II and IV spaces, $\mathscr{M}$ and $\mathscr{F}$ are both trivial, hence $P=W$ refines them. This is a strict refinement for Painlevé II and IV, as there $P=W$ has two steps, see [Sza21, Table 2.].

### v2

@@ 3.4. Painlevé IV (degenerate) case
With the above assumptions, the trace of $\theta$ vanishes, and its determinant reads as
@@
This follows from [ISS19a, Equation (70)]. Note that in that formula, $z_{2}=z^{-1}$, and we made use of the trivialization $\frac{\operatorname{d}\!z_{2}}{z_{2}}$ of $K(D)$ near $z=\infty$. Furthermore, the parameter $b_{-1}$ vanishes, and our free parameter $b$ can be identified with the parameter $t$ appearing in [ISS19a]. Finally, $w_{2}=\zeta$. ∎
@@
There exists a holomorphic $\mathbb{C}^{*}$-action on $\mathcal{M}^{PIV}$, equivariant with respect to the action $b\mapsto t^{3}b$ on $\mathcal{B}^{PIV}$. It acts on $\Omega_{I}^{PIV}$ with weight $2$.
@@
in particular $F_{\mathrm{min}}=F_{0}$. Therefore, subspaces of $H_{1}$ corresponds to flow within $\mathbb{C}P^{1}$’s, thus the weight decomposition of $F_{1},F_{2}$ and $F_{3}$ has $H_{-1},$ thus (due to 2.3) $H_{3}$ weight space:
(32) $T_{F_{1}}Y=H_{-1}\oplus H_{3},\quad T_{F_{2}}Y=H_{-1}\oplus H_{3},\quad T_{F_{3}}Y=H_{-1}\oplus H_{3},$
which goes outside of the core. As before, the $\mathbb{C}^{*}$-flowline corresponding to $H_{3}$ from one of these fixed points is the Hitchin section. Now, computation of the indices (11) gives:151515where, as before we denote $\mu_{\lambda+}:=\mu_{\lambda+\delta}$ for $01\gg\delta>$ 0
@@
The column $H^{*}(B_{1})$ is determined by 2.10, as the intersection form of $\mathcal{M}^{PIV}$ has a one-dimensional kernel, spanned by the sum of core components (labeled by $E_{1,2,3}$):
@@ 3.5. Painlevé VI case
the $\mathbb{C}^{*}$-fixed locus has to have a total cohomology rank of 6. The $\mathbb{C}P^{1}$ corresponding to the central vertex has 4 fixed points, hence it is fixed. Due to the smoothness of fixed loci, no other $\mathbb{C}P^{1}$s are fixed, hence they have another fixed point each, thus:
@@
and on the other hand $H^{0}(F_{i})[-2]=\langle[E_{i}]\rangle$ in $H^{2}(Y)$, where $E_{i}$ is the $\mathbb{C}P^{1}$ which $F_{i}$ belongs to. By 2.6, $1\in H^{0}(F_{0})$ enters the filtration at $\lambda=2,$ hence altogether we have:
(36) $\begin{split}\mathscr{F}_{1/2}\cong\mathbb{K}^{4}\subset\mathscr{F}_{1}=H^{2}(Y;\mathbb{K})\subset\mathscr{F}_{2}=H^{*}(Y;\mathbb{K})\\ \mathscr{F}_{\mathbb{B},1/2}=\langle[E_{1}],\dots,[E_{4}]\rangle\subset\mathscr{F}_{\mathbb{B},1}=H^{2}(Y)\subset\mathscr{F}_{2}=H^{*}(Y)\end{split}$
From the weight decompositions (34),(35), we get the outer torsion manifolds $Y_{2}\cong\mathbb{C}\sqcup\mathbb{C}\sqcup\mathbb{C}\sqcup\mathbb{C},$ thus $B_{1/2}\cong S^{1}\sqcup S^{1}\sqcup S^{1}\sqcup S^{1}.$ Their shifts, computed via (15), give the columns of the $E_{1}$-page of the spectral sequence, shown on Figure 4. The column $H^{*}(B_{1})$ is determined by 2.10, as the intersection form of $\mathcal{M}^{PVI}$ has a one-dimensional kernel, spanned by the weighted sum $2[F_{0}]+[E_{i}]$ of the core components. Like in the previous example, the spectral sequence alone does not completely determine the filtration (36), as we could not tell which of the columns $B_{1},B_{3/2}$ and $B_{2}$ annihilates $1\in H^{0}(\mathcal{M}^{PIV})$.
@@ 3.6. Remaining Painlevé cases
3.7. Further Higgs moduli of $\dim_{\mathbb{C}}=2$
@@
together with $\pi$ induces Hitchin fibration
(38) $h_{\Gamma}:\mathcal{M}_{\Gamma}\rightarrow(T^{*}E)/\Gamma\rightarrow\mathbb{C}/\Gamma=\mathcal{B}_{\Gamma}\cong\mathbb{C}.$
@@
[Zha17, Proposition 5.4] The central fiber of (38) is
@@
The fiber-contraction on $T^{*}E$ that makes (37) equivariant (with weight-1 action on $\mathbb{C}$) induces a $\mathbb{C}^{*}$-action on $(T^{*}E)/\Gamma,$ which lifts on $\mathcal{M}_{\Gamma}$ as it is a crepant resolution of a “nice” isolated singularity (one can use e.g. [MR23, Proposition 3.6.]). This altogether makes (38) equivariant. For completeness, we describe this action in the Higgs bundle-theoretic way:
@@
$\zeta^{4}+b\frac{(\operatorname{d}\!z)^{\otimes 4}}{z^{3}(z-1)^{3}}=0.$
@@
$\zeta^{6}+b\frac{(\operatorname{d}\!z)^{\otimes 6}}{z^{5}(z-1)^{4}}=0,$
@@
The Floer-theoretic filtration $\mathscr{F}$ with respect to this $\mathbb{C}^{*}$-action on $\mathcal{M}_{\Gamma}$ is studied in [RŽ23b], where it is compared with the perverse filtration of Hitchin fibration $h_{\Gamma}:\mathcal{M}_{\Gamma}\rightarrow\mathcal{B}_{\Gamma}$ (the famous “$P=W$” filtration of Higgs moduli spaces). In addition, there is an interesting correlation with the associated root system of graph $Q_{\Gamma}.$ Being an affine Dynkin graph, it has the imaginary root, the positive $\mathbb{Z}$-generator of the radical of the symmetric bilinear form with associated Cartan matrix $C_{Q_{\Gamma}}.$ One can label the vertices $i\in Q_{\Gamma}^{0}$ by the corresponding entries $n_{i}$ of the imaginary root, see Figure 5.
@@
Given two filtrations $\{F_{i}\}_{i}$ and $\{G_{j}\}_{j}$ of a vector space $V,$ we say that $F$ refines $G$ if for each $j$ there is some $i$ such that $G_{j}=F_{i}$.
@@
3.8. Classification of 2-dimensional $\mathbb{C}^{*}$-equivariant Higgs moduli spaces
Proposition 3.15.
Let $\mathcal{M}$ be a moduli space of Higgs bundles over some curve $X$. Assume that
  1. (1)
    $\mathcal{M}$ is of complex dimension $2$;
  2. (2)
    there exists a smooth Hitchin map $h\colon\mathcal{M}\to\mathbb{C}$ turning $\mathcal{M}$ into a smooth elliptic fibration;
  3. (3)
    $\mathbb{C}^{*}$ acts holomorphically on $\mathcal{M}$ and $h$ is equivariant for some contracting action on $\mathbb{C}.$
Then, $\mathcal{M}$ is biregular to one of the spaces $\mathcal{M}^{PX}$ or $\mathcal{M}_{\Gamma}$ described in Sections 3.13.7. Moreover, $h$ agrees with the corresponding Hitchin map of $\mathcal{M}^{PX}$ and $\mathcal{M}_{\Gamma}$ too.
Proof.
It follows from assumption (2) that $\mathcal{M}$ is a Zariski open subset of the rational elliptic surface $\operatorname{Bl}_{9}\mathbb{C}P^{2}$ ($\mathbb{C}P^{2}$ blown up $9$ times). The Hitchin map is algebraic, therefore it extends to an elliptic fibration
$\tilde{h}\colon\operatorname{Bl}_{9}\mathbb{C}P^{2}\to\mathbb{C}P^{1}.$
From assumption (3) and smoothness of $h$ we infer that $\tilde{h}$ may have at most two singular fibers, namely the fibers over $0,\infty\in\mathbb{C}P^{1}$.
If $h$ admits no singular fiber, then $\mathcal{M}$ is diffeomorphic to $T^{*}E$ for an elliptic curve $E$, i.e. $\mathcal{M}_{\Gamma}$ for $\Gamma=0$.
Assume that $h$ admits exactly one singular fiber, and $\tilde{h}$ has exactly two singular fibers. Then, from the structure of the monodromy of the Kodaira fibers, it is well-known that the only possibilities for couples of singular fibers over $\operatorname{Bl}_{9}\mathbb{C}P^{2}$ are $I_{0}^{*}+I_{0}^{*}$, $\widetilde{E}_{6}+IV$, $\widetilde{E}_{7}+III$ or $\widetilde{E}_{8}+II$.
The case $I_{0}^{*}+I_{0}^{*}$ is symmetric. In the other cases, the fiber over 00 can be either of the two singular fibers. In Sections 3.13.7 we have found all these possibilities. ∎
Remark 3.16.
We would like to emphasize that there exist different moduli problems from the ones detailed here that have diffeomorphic moduli spaces. Namely, in [Boa12, Section 11] different ”readings” of a quiver are introduced and studied. For instance, in [Boa12, Section 11.3], infinitely many Lax representations of $PIV$ are described. It would be interesting to work out in detail the Hitchin map for such other spaces explicitly. We leave this question for further study. The main challenge is that the spectral correspondence produces many-sheeted covers of $\mathbb{C}P^{1}$, which then get birationally transformed to high degree (and thus high genus) curves in $\mathbb{C}P^{2}$, whose Jacobi varieties are correspondingly of high dimension. Therefore, in general it is not clear to the authors how they give rise to elliptic surfaces. However, for the rank $3$ Lax pairs of the Painlevé equations, in [ES25] it is proven that Fourier–Laplace transformation establishes symplectic isomorphisms with the Hitchin moduli spaces treated here.
@@ 4. Three filtrations compared
The imaginary root labels mentioned above have another geometric meaning. Recall from  3.11 that $\mathrm{Core}(\mathcal{M}_{\Gamma})=h_{\Gamma}^{-1}(0)=\cup E_{i}$ has irreducible components $E_{i}$ labeled by vertices $i$ of $Q_{\Gamma}.$ Moreover, its canonical scheme structure comes with multiplicities $m_{i}$ of curves $E_{i}$
@@
with the labelings $n_{i}$ of the imaginary root of $Q_{\Gamma}.$ This is a consequence of
@@
    Self-intersection of $h_{\Gamma}^{-1}(0)$ (with multiplicities) in the sublattice spanned by $\mathrm{Irr}(\mathrm{Core}(\mathcal{M}_{\Gamma}))$ is equal to zero, as a fiber class.
@@
    By definition, the imaginary root is the positive integer generator of the radical of the symmetric bilinear form with given Cartan matrix, which equals minus the intersection matrix of the elements $E_{i}$ of $\mathrm{Irr}(\mathrm{Core}(\mathcal{M}))$.
@@
(40) $\mathrm{rk}\,\mathscr{F}_{k/|\Gamma|}=\#\{E_{i}\in\mathrm{Irr}(\mathrm{Core}(\mathcal{M}_{\Gamma}))\mid m_{i}\leq k\}$
@@
Corollary 4.2.
For any 2-dimensional parabolic Higgs bundle $\mathcal{M}_{\Gamma},$ we have
@@
Proof.
Notice by (39) that the maximal multiplicity $m$ of the irreducible core components corresponds to the maximal labeled number of the associated imaginary root. Looking at Figure 5, we see that this maximal root label is precisely $|\Gamma|$, as for $\widetilde{D}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8}$, the corresponding groups $\Gamma$ are (as mentioned before) $\mathbb{Z}/2,\mathbb{Z}/3,\mathbb{Z}/4,\mathbb{Z}/6,$ respectively. For $\widetilde{A}_{0},$ the multiplicity of the (only) component is $1,$ so again equal to $|\Gamma|$, as $\Gamma=0$ in that case. Thus, $m=|\Gamma|$, and by (40)
$\mathrm{rk}\,\mathscr{F}_{k/\Gamma}=\#\{E_{i}\mid m_{i}\leq k\}=\mathrm{rk}\,\langle[E_{i}]\mid m_{i}/|\Gamma|\leq k/|\Gamma|\rangle=\langle[E_{i}]\mid m_{i}/m\leq k/|\Gamma|\rangle=\mathrm{rk}\,\mathscr{M}_{k/\Gamma},$
and both filtrations only jump in rank for $\lambda=k/\Gamma,$ hence (41) is proved. ∎
Following ideas from [RŽ23a], one expects that these filtrations actually coincide, not only rank-wise:
@@
Let us be given any Higgs moduli space $\mathcal{M}$ satisfying the conditions of Proposition 3.15. Then, its multiplicity and Floer-theoretic filtration coincide rank-wise
$(\forall\lambda)\ \mathrm{rk}\,\mathscr{F}_{\lambda}=\mathrm{rk}\,\mathscr{M}_{\lambda}.$
For moduli spaces of parabolic Higgs bundles, they both refine the $P=W$ filtration, whereas for Painlevé spaces, $P=W$ refines the former two filtrations. Thus, in the intersection, i.e. for Painlevé VI space, these three filtrations coincide.
@@
By Proposition 3.15, we only need to deal with the cases covered in Sections 3.13.7. By (41), the multiplicity and Floer-theoretic filtration coincide rank-wise for parabolic spaces $\mathcal{M}_{\Gamma}.$ For Painlevé spaces $\mathcal{M}^{PX},\ X\in\{I,II,IV\}$, both filtrations are trivial, so they coincide. The Floer-theoretic filtration $\mathscr{F}=\mathscr{F}_{\mathbb{B}}(H^{2}(\mathcal{M}^{PX}))$ is trivial due to (22),(29),(33), respectively. The multiplicity filtration is trivial as the multiplicities of core components $E_{i}$ are all $m_{i}=1$ in these cases. The reasoning for this is similar to that for (39), as the self-intersection of $\sum_{i}E_{i}$ is zero. The latter has already been proved at the ends of their respective Sections 3.1, 3.3 and 3.4.
For the parabolic moduli spaces $\mathcal{M}_{\Gamma}$, the multiplicity filtration $\mathscr{M}$ refines the $P=W$, as for $\Gamma\neq 0,$ the latter has two steps (see [Zha17, Proposition 5.4]):
@@
For Painlevé I, II, and IV spaces, $\mathscr{M}$ and $\mathscr{F}$ are both trivial, hence $P=W$ refines them. This is a strict refinement for Painlevé II and IV, as there $P=W$ has two steps, see [Sza21, Table 2.].